# A Vision for Trustworthy, Fair, and Efficient Socio-Technical Control using Karma Economies

### v1

@@ 3.2 The fairness and efficiency of Maximum Long-run Nash Welfare (MLNW)
Figure 6 illustrates the MLNW solution (Figure 6(b)) for allocating highway priority lanes over the morning rush hour, as specified in Example 4. We additionally consider that the users are uniformly grouped in one of the four urgency processes shown in Figure 6(a), which all have the same average urgency but different dynamics. To highlight the fairness and efficiency of MLNW, it is contrasted to two other solutions: Figure 6(c) shows the outcomes if Nash welfare is maximized separately for each day in a single-shot manner; and Figure 6(d) shows the utilitarian solution, which assumes that the urgency is interpersonally comparable and allocates the most desirable resources to the highest urgency irrespective of the underlying urgency process. Contrasting first to the single-shot maximum Nash welfare solution, it allocates all users their fair share of each departure interval irrespective of urgency, due to its scale invariance property, cf. Figure 6(c)–left. While this still leads to strong Pareto improvements in the long-run expected average rewards over the non-controlled benchmark, cf. Figure 6(c)–right, it is pronouncedly dominated by MLNW that properly accounts for the dynamics, cf. Figure 6(b)–right. It can be seen that MLNW extracts possible efficiency gains from the urgency dynamics in an interpersonally non-comparable, and thereby fairness preserving, manner. Namely, as revealed by Figure 6(b)–left, different users ‘trade’ their fair shares of different departure intervals in a mutually beneficial manner. For example, users who experience urgency $u_{\textup{h}}=9$ one eighth of the time (displayed green) give up the majority of their fair shares of most departure intervals in exchange of a larger share of the most desirable intervals $j^{*}$ and $j^{*}-1$. On the other hand, users who experience a constant urgency $u=2$ (displayed red), while they give up some of their fair shares of $j^{*}$, $j^{*}-1$, $j^{*}-2$, gain a larger share of all other intervals.
Contrasting second to the utilitarian solution demonstrates the consequences of naively assuming that urgency is interpersonally comparable. In this case, users who experience a constant urgency $u=2$ (displayed red) are never allocated to the priority lanes, because there are always sufficiently many users with ‘higher’ urgency to fill those lanes, cf. Figure 6(d)–left. As a consequence, those users do not realize any benefits over the non-controlled benchmark, cf. Figure 6(d)–right. If one interprets the urgency process as how often a user is late to work, those who are always disciplined in adhering to schedules are punished under utilitarianism. Naturally, in the absence of money or a fair instrument to make interpersonal comparisons, one would expect these users to over-report their urgency to realize benefits under a utilitarian scheme.
@@ 4 Dynamic Resource Allocation with Karma Economies
In this section, we formally introduce karma economies for dynamic resource allocation (Section 4.1) and their game-theoretical equilibrium notion of Karma Equilibrium (KE) (Section 4.2). The main result of this section, which establishes that karma economies implement MLNW in a decentralized and trustworthy (Theorem 5), follows in Section 4.3. Moreover, Section 4.4 discusses a complementary modelling tool for karma economies that produces finer-grained equilibrium predictions, while Section 4.5 contrasts karma economies to related mechanisms in the economics and computer science literature, highlighting the ‘closed-loop’ nature of karma.
@@ 4.2 Game-theoretical solution concept: Karma Equilibrium (KE)
{mdframed}
[backgroundcolor=red!10] Karma User Problem: $\max_{\bm{\chi_{i}}}$ $\bar{r}_{i}-\bar{r}^{0}_{i}$ (KU) subject to $\sum_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\sum_{j\in{\mathcal{M}}}\chi_{i,j}(u_{i})\,b^{\star}_{j}\leq\frac{w_{i}}{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}\sum_{j\in{\mathcal{M}}}b^{\star}_{j}\,c_{j},$ $0\leq\chi_{i,j}(u_{i})\leq 1,\;\forall j\in{\mathcal{M}},u_{i}\in{\mathcal{U}}_{i}.$
@@
A Karma Equilibrium (KE) is a pair $\left(\bm{\chi^{\star}},\bm{b^{\star}}\right)$ of long-run allocations $\bm{\chi^{\star}}=\left(\bm{\chi^{\star}_{i}}\right)_{i\in{\mathcal{N}}}$ and stationary bids $\bm{b^{\star}}=\left(b^{\star}_{j}\right)_{j\in{\mathcal{M}}}$ that satisfy:
  1. 1.
    Individual Optimality: For all users $i\in{\mathcal{N}}$, $\bm{\chi^{\star}_{i}}$ solves Problem (KU) given $\bm{b^{\star}}$;
  2. 2.
    Resource Clearing: For all resources $j\in{\mathcal{M}}$, $\sum\limits_{i\in{\mathcal{N}}}\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\,\chi^{\star}_{i,j}(u_{i})\leq c_{j}$, and $b^{\star}_{j}>0$ implies that $\sum\limits_{i\in{\mathcal{N}}}\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\,\chi^{\star}_{i,j}(u_{i})=c_{j}$;
  3. 3.
    Budget Balance: For all users $i\in{\mathcal{N}}$, $\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\sum\limits_{j\in{\mathcal{M}}}\chi^{\star}_{i,j}(u_{i})\,b^{\star}_{j}=\cfrac{w_{i}}{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}\sum\limits_{j\in{\mathcal{M}}}b^{\star}_{j}\,c_{j}$.
@@
Let Assumption 1 hold. Then KE $(\bm{\chi^{\star}},\bm{b^{\star}},\bm{\eta^{\star}})$ satisfy, for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$:
  1. 4.1
    $\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}>0$;
  2. 4.2
@@
    If two users $i$, $i^{\prime}$ are symmetric up to scales $\alpha_{i},\alpha_{i^{\prime}}>0$, then $\cfrac{\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}}{\alpha_{i}}=\cfrac{\left(\bar{r}_{i^{\prime}}-\bar{r}^{0}_{i^{\prime}}\right)^{\star}}{\alpha_{i^{\prime}}}$;
  3. 4.3
    $\chi^{\star}_{i,j}(u_{i})>0\Rightarrow\chi^{\star}_{i,j}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$;
  4. 4.4
    $b^{\star}_{j}>0$, and all resources are allocated at capacity;
  5. 4.5
@@
    $u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)\leq\cfrac{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}{w_{i}\sum_{j^{\prime}\in{\mathcal{M}}}b^{\star}_{j^{\prime}}\,c_{j^{\prime}}}\left(\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}-\sum\limits_{u^{\prime}_{i}\in{\mathcal{U}}_{i}}\sum\limits_{j^{\prime}\in{\mathcal{M}}}\eta^{\star}_{i,j^{\prime}}(u^{\prime}_{i})\right)b^{\star}_{j}+\cfrac{\eta^{\star}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}$;
  6. 4.6
    $\chi^{\star}_{i,j}(u_{i})>0\Rightarrow u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)$
    $=\cfrac{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}{w_{i}\sum_{j^{\prime}\in{\mathcal{M}}}b^{\star}_{j^{\prime}}\,c_{j^{\prime}}}\left(\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}-\sum\limits_{u^{\prime}_{i}\in{\mathcal{U}}_{i}}\sum\limits_{j^{\prime}\in{\mathcal{M}}}\eta^{\star}_{i,j^{\prime}}(u^{\prime}_{i})\right)b^{\star}_{j}+\cfrac{\eta^{\star}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}$;
  7. 4.7
@@
    If there is only one resource =${\mathcal{M}}=\{1\}$, then $\sum_{u_{i}\in{\mathcal{U}}_{i}}\sigma(u_{i})\chi^{\star}_{i,1}(u_{i})=\frac{w_{i}}{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}\,c_{1}$.
@@
Moreover, if the resources are mutually exclusive, then Propositions 4.14.2 hold, Propositions 4.54.6 hold with non-resource specific multipliers $\eta^{\star}_{i,j}(u_{i})=\eta^{\star}_{i}(u_{i})$, and it holds for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$:
  1. 4.3
    $\chi^{\star}_{i,j}(u_{i})>0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{\star}_{i,j^{\prime}}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$;
  2. 4.4
    $b^{\star}_{j}=0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{\star}_{i,j^{\prime}}(u_{i})=1$, for all $i\in{\mathcal{N}}_{j}$, $u_{i}\in{\mathcal{U}}_{i}$.
@@
The proof of Proposition 4 is included in A.2. The KE mirrors the properties of MLNW, cf. Proposition 3, with the equilibrium clearing bids $\bm{b^{\star}}$ taking a similar role as the shadow prices $\bm{\lambda^{*}}$ under MLNW. However, a connection between both solution concepts is not immediate, as comparing Propositions 4.54.6 to Propositions 3.53.6 reveals: constraints $\chi_{i,j}(u_{i})\leq 1$ and the associated optimal multipliers $\eta^{\star}_{i,j}(u_{i})$ affect the solution differently in both concepts. Moreover, in single resource settings, the KE guarantees that each user’s long-run expected average allocation equals its fair share, cf. Proposition 4.7, which in general is not guaranteed to hold under NLNW. In what follows, we establish conditions under which both solution concepts coincide.
@@ 4.3 Trustworthy MLNW implementation with karma economies
A dynamic resource allocation problem ${\mathfrak{D}}$ is said to be Nash-balanced if it admits a MLNW solution $\left(\bm{\chi^{*}},\bm{\lambda^{*}},\bm{\eta^{*}}\right)$ satisfying, for all users $i,i^{\prime}\in{\mathcal{N}}$
$\frac{\sum_{j\in{\mathcal{M}}}\sum_{u_{i}\in{\mathcal{U}}_{i}}\eta^{*}_{i,j}(u_{i})}{w_{i}}=\frac{\sum_{j\in{\mathcal{M}}}\sum_{u_{i^{\prime}}\in{\mathcal{U}}_{i^{\prime}}}\eta^{*}_{i^{\prime},j}(u_{i^{\prime}})}{w_{i^{\prime}}}=C.$ (NB)
@@
While it remains an open problem to fully characterize Nash-balanced dynamic resource allocation problems, we discuss when condition (NB) is expected to hold and argue that it is not severely restrictive. A natural setting in which condition (NB) is satisfied is when constraints $\chi^{*}_{i,j}\leq 1$ do not bind for any $i\in{\mathcal{U}}_{i}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$, and thus $C=\frac{\sum_{u_{i}\in{\mathcal{U}}_{i}}\sum_{j\in{\mathcal{M}}}\eta^{*}_{i,j}(u_{i})}{w_{i}}=0$ for all $i\in{\mathcal{N}}$. By Proposition 3.3, this occurs when there is resource scarcity leading to MLNW allocations to users in their respective highest urgency only. Notice that this does not imply that the frequency and severity of the highest urgency must be the same for all users, or that the users end up with equal allocations: in fact, the afore-discussed example problem of allocating highway priority lanes over the morning rush hour (Example 5) is Nash-balanced, with $\eta^{*}_{i,j}(u_{i})=\eta^{*}_{i}(u_{i})=0$ for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$, and the MLNW solution portrayed in Figure 6(b) is indeed a KE of this problem.
Moreover, there exist Nash-balanced problems for which $C>0$ and thus constraints $\chi^{*}_{i,j}(u_{i})\leq 1$ bind for each user $i\in{\mathcal{N}}$ at some resource $j\in{\mathcal{M}}$ and urgency $u_{i}\in{\mathcal{U}}_{i}$. Figure 7 shows an example with a single resource =${\mathcal{M}}=\{1\}$ of capacity $c_{1}=1$, equal access rights, and users belonging to two pairs of urgency processes: one pair leads to a Nash-balanced problem with $C>0$ (Figure 7(a)) and the other does not lead to a Nash-balanced problem (Figure 7(b)). By Proposition 4.7, the KE allocates users of both processes the fair share of $0.5$ in both problems, while this occurs under MLNW only in the Nash-balanced problem, cf. Figure 7(a). In the non-Nash-balanced problem, it is optimal under MLNW to allocate users of one of the urgency processes more than fair share, cf. Figure 7(b). Intuitively, one would not expect this outcome to be implementable in a game-theoretic equilibrium, as users of the urgency process that receives less than fair share would increase their allocation if they over-report the time spent in high urgency. In mathematical terms, those users are penalized disproportionally for satisfying constraints $\chi^{*}_{i,j}(u_{i})\leq 1$ under MLNW.

### v2

@@ 3.2 The fairness and efficiency of Maximum Long-run Nash Welfare (MLNW)
Figure 6 illustrates the MLNW solution (Figure 6(b)) for allocating highway priority lanes over the morning rush hour, as specified in Example 4. We additionally consider that the users are uniformly grouped in one of the four urgency processes shown in Figure 6(a), which all have the same average urgency but different dynamics [^15] Optimization problem (MLNW) is solved using a representative user of each urgency process, rather than all 9000 users, leveraging the symmetry property of Proposition 3.2.. To highlight the fairness and efficiency of MLNW, it is contrasted to two other solutions: Figure 6(c) shows the outcomes if Nash welfare is maximized separately for each day in a single-shot manner; and Figure 6(d) shows the utilitarian solution, which assumes that the urgency is interpersonally comparable and allocates the most desirable resources to the highest urgency irrespective of the underlying urgency process. Contrasting first to the single-shot maximum Nash welfare solution, it allocates all users their fair share of each departure interval irrespective of urgency, due to its scale invariance property, cf. Figure 6(c)–left. While this still leads to strong Pareto improvements in the long-run expected average rewards over the non-controlled benchmark, cf. Figure 6(c)–right, it is pronouncedly dominated by MLNW that properly accounts for the dynamics, cf. Figure 6(b)–right. It can be seen that MLNW extracts possible efficiency gains from the urgency dynamics in an interpersonally non-comparable, and thereby fairness preserving, manner. Namely, as revealed by Figure 6(b)–left, different users ‘trade’ their fair shares of different departure intervals in a mutually beneficial manner. For example, users who experience urgency $u_{\textup{h}}=9$ one eighth of the time (displayed green) give up the majority of their fair shares of most departure intervals in exchange of a larger share of the most desirable intervals $j^{*}$ and $j^{*}-1$. On the other hand, users who experience a constant urgency $u=2$ (displayed red), while they give up some of their fair shares of $j^{*}$, $j^{*}-1$, $j^{*}-2$, gain a larger share of all other intervals. Contrasting second to the utilitarian solution demonstrates the consequences of naively assuming that urgency is interpersonally comparable. In this case, users who experience a constant urgency $u=2$ (displayed red) are never allocated to the priority lanes, because there are always sufficiently many users with ‘higher’ urgency to fill those lanes, cf. Figure 6(d)–left. As a consequence, those users do not realize any benefits over the non-controlled benchmark, cf. Figure 6(d)–right. If one interprets the urgency process as how often a user is late to work, those who are always disciplined in adhering to schedules are punished under utilitarianism. Naturally, in the absence of money or a fair instrument to make interpersonal comparisons, one would expect these users to over-report their urgency to realize benefits under a utilitarian scheme.
@@ 4 Dynamic Resource Allocation with Karma Economies
In this section, we formally introduce karma economies for dynamic resource allocation (Section 4.1) and their game-theoretical equilibrium notion of Karma Equilibrium (KE) (Section 4.2). The main result of this section, which establishes that karma economies implement MLNW in a decentralized and trustworthy manner (Theorem 5), follows in Section 4.3. Moreover, Section 4.4 discusses a complementary modelling tool for karma economies that produces finer-grained equilibrium predictions, while Section 4.5 contrasts karma economies to related mechanisms in the economics and computer science literature, highlighting the ‘closed-loop’ nature of karma.
@@ 4.2 Game-theoretical solution concept: Karma Equilibrium (KE)
Karma User Problem: $\max_{\bm{\chi_{i}}}$ $\bar{r}_{i}-\bar{r}^{0}_{i}$ (KU) subject to $\sum_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\sum_{j\in{\mathcal{M}}}\chi_{i,j}(u_{i})\,b^{\star}_{j}\leq\frac{w_{i}}{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}\sum_{j\in{\mathcal{M}}}b^{\star}_{j}\,c_{j},$ $0\leq\chi_{i,j}(u_{i})\leq 1,\;\forall j\in{\mathcal{M}},u_{i}\in{\mathcal{U}}_{i}.$
@@
A Karma Equilibrium (KE) is a pair $\left(\bm{\chi^{\star}},\bm{b^{\star}}\right)$ of long-run allocations $\bm{\chi^{\star}}=\left(\bm{\chi^{\star}_{i}}\right)_{i\in{\mathcal{N}}}$ and stationary bids $\bm{b^{\star}}=\left(b^{\star}_{j}\right)_{j\in{\mathcal{M}}}$ that satisfy: 1. Individual Optimality: For all users $i\in{\mathcal{N}}$, $\bm{\chi^{\star}_{i}}$ solves Problem (KU) given $\bm{b^{\star}}$; 2. Resource Clearing: For all resources $j\in{\mathcal{M}}$, $\sum\limits_{i\in{\mathcal{N}}}\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\,\chi^{\star}_{i,j}(u_{i})\leq c_{j}$, and $b^{\star}_{j}>0$ implies that $\sum\limits_{i\in{\mathcal{N}}}\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\,\chi^{\star}_{i,j}(u_{i})=c_{j}$; 3. Budget Balance: For all users $i\in{\mathcal{N}}$, $\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\sum\limits_{j\in{\mathcal{M}}}\chi^{\star}_{i,j}(u_{i})\,b^{\star}_{j}=\cfrac{w_{i}}{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}\sum\limits_{j\in{\mathcal{M}}}b^{\star}_{j}\,c_{j}$.
@@
Let Assumption 1 hold. Then KE $(\bm{\chi^{\star}},\bm{b^{\star}},\bm{\eta^{\star}})$ satisfy, for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$: 4.1 $\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}>0$; 4.2 If two users $i$, $i^{\prime}$ are symmetric up to scales $\alpha_{i},\alpha_{i^{\prime}}>0$, then $\cfrac{\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}}{\alpha_{i}}=\cfrac{\left(\bar{r}_{i^{\prime}}-\bar{r}^{0}_{i^{\prime}}\right)^{\star}}{\alpha_{i^{\prime}}}$; 4.3 $\chi^{\star}_{i,j}(u_{i})>0\Rightarrow\chi^{\star}_{i,j}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$; 4.4 $b^{\star}_{j}>0$, and all resources are allocated at capacity; 4.5 $u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)\leq\cfrac{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}{w_{i}\sum_{j^{\prime}\in{\mathcal{M}}}b^{\star}_{j^{\prime}}\,c_{j^{\prime}}}\left(\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}-\sum\limits_{u^{\prime}_{i}\in{\mathcal{U}}_{i}}\sum\limits_{j^{\prime}\in{\mathcal{M}}}\eta^{\star}_{i,j^{\prime}}(u^{\prime}_{i})\right)b^{\star}_{j}+\cfrac{\eta^{\star}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}$; 4.6 $\chi^{\star}_{i,j}(u_{i})>0\Rightarrow u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)$
@@
$\cfrac{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}{w_{i}\sum_{j^{\prime}\in{\mathcal{M}}}b^{\star}_{j^{\prime}}\,c_{j^{\prime}}}\left(\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}-\sum\limits_{u^{\prime}_{i}\in{\mathcal{U}}_{i}}\sum\limits_{j^{\prime}\in{\mathcal{M}}}\eta^{\star}_{i,j^{\prime}}(u^{\prime}_{i})\right)b^{\star}_{j}+\cfrac{\eta^{\star}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}$;
@@
4.7 If there is only one resource ${\mathcal{M}}=\{1\}$, then $\sum_{u_{i}\in{\mathcal{U}}_{i}}\sigma(u_{i})\chi^{\star}_{i,1}(u_{i})=\frac{w_{i}}{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}\,c_{1}$.
@@
Moreover, if the resources are mutually exclusive, then Propositions 4.14.2 hold, Propositions 4.54.6 hold with non-resource specific multipliers $\eta^{\star}_{i,j}(u_{i})=\eta^{\star}_{i}(u_{i})$, and it holds for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$: 4.3 $\chi^{\star}_{i,j}(u_{i})>0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{\star}_{i,j^{\prime}}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$; 4.4 $b^{\star}_{j}=0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{\star}_{i,j^{\prime}}(u_{i})=1$, for all $i\in{\mathcal{N}}_{j}$, $u_{i}\in{\mathcal{U}}_{i}$.
@@
The proof of Proposition 4 is included in A.2. The KE mirrors the properties of MLNW, cf. Proposition 3, with the equilibrium clearing bids $\bm{b^{\star}}$ taking a similar role as the shadow prices $\bm{\lambda^{*}}$ under MLNW. However, a connection between both solution concepts is not immediate, as comparing Propositions 4.54.6 to Propositions 3.53.6 reveals: constraints $\chi_{i,j}(u_{i})\leq 1$ and the associated optimal multipliers $\eta^{\star}_{i,j}(u_{i})$ affect the solution differently in both concepts. Moreover, in single resource settings, the KE guarantees that each user’s long-run expected average allocation equals its fair share, cf. Proposition 4.7, which in general is not guaranteed to hold under MLNW. In what follows, we establish conditions under which both solution concepts coincide.
@@ 4.3 Trustworthy MLNW implementation with karma economies
A dynamic resource allocation problem $\mathfrak{D}$} is said to be Nash-balanced if it admits a MLNW solution $\left(\bm{\chi^{*}},\bm{\lambda^{*}},\bm{\eta^{*}}\right)$ satisfying, for all users $i,i^{\prime}\in{\mathcal{N}}$ $\frac{\sum_{j\in{\mathcal{M}}}\sum_{u_{i}\in{\mathcal{U}}_{i}}\eta^{*}_{i,j}(u_{i})}{w_{i}}=\frac{\sum_{j\in{\mathcal{M}}}\sum_{u_{i^{\prime}}\in{\mathcal{U}}_{i^{\prime}}}\eta^{*}_{i^{\prime},j}(u_{i^{\prime}})}{w_{i^{\prime}}}=C.$ (NB)
@@
While it remains an open problem to fully characterize Nash-balanced dynamic resource allocation problems, we discuss when condition (NB) is expected to hold and argue that it is not severely restrictive. A natural setting in which condition (NB) is satisfied is when constraints $\chi^{*}_{i,j}\leq 1$ do not bind for any $i\in{\mathcal{U}}_{i}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$, and thus $C=\frac{\sum_{u_{i}\in{\mathcal{U}}_{i}}\sum_{j\in{\mathcal{M}}}\eta^{*}_{i,j}(u_{i})}{w_{i}}=0$ for all $i\in{\mathcal{N}}$. By Proposition 3.3, this occurs when there is resource scarcity leading to MLNW allocations to users in their respective highest urgency only. Notice that this does not imply that the frequency and severity of the highest urgency must be the same for all users, or that the users end up with equal allocations: in fact, the afore-discussed example problem of allocating highway priority lanes over the morning rush hour (Example 5) is Nash-balanced, with $\eta^{*}_{i,j}(u_{i})=\eta^{*}_{i}(u_{i})=0$ for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$, and the MLNW solution portrayed in Figure 6(b) is indeed a KE of this problem. Moreover, there exist Nash-balanced problems for which $C>0$ and thus constraints $\chi^{*}_{i,j}(u_{i})\leq 1$ bind for each user $i\in{\mathcal{N}}$ at some resource $j\in{\mathcal{M}}$ and urgency $u_{i}\in{\mathcal{U}}_{i}$. Figure 7 shows an example with a single resource ${\mathcal{M}}=\{1\}$ of capacity $c_{1}=1$, equal access rights, and users belonging to two pairs of urgency processes: one pair leads to a Nash-balanced problem with $C>0$ (Figure 7(a)) and the other does not lead to a Nash-balanced problem (Figure 7(b)). By Proposition 4.7, the KE allocates users of both processes the fair share of $0.5$ in both problems, while this occurs under MLNW only in the Nash-balanced problem, cf. Figure 7(a). In the non-Nash-balanced problem, it is optimal under MLNW to allocate users of one of the urgency processes more than fair share, cf. Figure 7(b). Intuitively, one would not expect this outcome to be implementable in a game-theoretic equilibrium, as users of the urgency process that receives less than fair share would increase their allocation if they over-report the time spent in high urgency. In mathematical terms, those users are penalized disproportionally for satisfying constraints $\chi^{*}_{i,j}(u_{i})\leq 1$ under MLNW.